# Near-field Beamforming for Extremely Large-scale MIMO Based on Unsupervised Deep Learning

### v1

@@ IV-B Experimental Result
Fig. 12 illustrates the average achievable rates by various schemes under different number of BS antennas $N_{t}$. In this experiment, The SNR is set to 10 dB, and other simulation parameters remained consistent with those in Table II. From the observation results, with the increase of the number of antennas, all schemes face a trend of performance degradation. Remarkably, our proposed scheme consistently exhibits superior performance compared to all comparison schemes. With the growth in the number of antennas, the performance of the proposed model improves more significantly than that of the comparison model. The escalating number of antennas makes it more difficult for traditional schemes to deal with intricate beam codebooks. This further substantiates the exceptional robustness and resilience of our proposed scheme under diverse conditions. The consistent outperformance of our model in the face of increasing antenna numbers underscores its adaptability and suitability for evolving communication scenarios.
V Conclusions
In this paper, we proposed a near-feild beamforming scheme based on unsupervised deep learning. By skillfully designing the padding and kernel size of the convolutional neural network, we effectively extract the features of the complex channel state information. During the network training process, we utilize the negative value of the total achievable rate in the multi-user network as the loss function and ensure the amplitude constraint of the beamforming vector through a customized Tanh layer. Our solution doesn’t rely on a predefined codebook and only requires the estimated channel state information as input to obtain the optimal beamforming vector. Simulation results demonstrate the competitive performance of the proposed scheme.

### v2

@@ V-B Experimental Result
Fig. 12 illustrates the average achievable rate under different numbers of BS antennas $N$. We can see that the performance of each scheme tends to decline with the increase in the number of antennas. Remarkably, our proposed scheme consistently exhibits superior compared to all comparison schemes. With the growth in the number of antennas, the performance of the proposed model improves more significantly than the compared model. The escalating number of antennas makes it more difficult for traditional schemes to deal with intricate beam codebooks. The consistent outperformance of our model in the face of increasing antenna numbers underscores its adaptability and suitability for evolving communication scenarios.
VI Conclusions
In this paper, we proposed a near-field beamforming scheme based on deep learning. By strategically designing the padding and kernel size of the convolutional neural network, we effectively extract the features of the complex CSI. During the network training process, we utilize the negative value of the total achievable rate in the multi-user network as the loss function and ensure the amplitude constraint of the beamforming vector through a customized Tanh layer. Our solution doesn’t rely on predefined codebooks and only requires the estimated CSI as input to obtain the optimal beamforming vector. Simulation results demonstrate the competitive performance of the proposed scheme.